OpenStudy (anonymous):

Concentrated sulfuric acid H2SO4 is 18.4M. What is concentration of a solution made by diluting 25.0 mL concentrated sulfuric acid with water to a final volume of 5.00 x 10^2 mL. I don't know how to set this up with dimensional analysis!

OpenStudy (callisto):

no. of mole of H2SO4 extracted for dilution = 18.4 X (25.0/1000) = 0.4625mole concentration required = 0.4625 / [(5X10^2)/1000] = 0.4625/0.5 = 0.925M